OpenStudy (anonymous):

Identify the correct formula for the following sequence. 14, 19, 24, 29, 34, ... a^n = 5n + 9 a^n = 45n a^n = 9n + 5 a^n = 5n - 9

OpenStudy (anonymous):

a^n = \[a _{n}\]

OpenStudy (jdoe0001):

hmm any ideas on the "common difference"?

OpenStudy (michele_laino):

I think: \[a _{n}=5n+9\]

OpenStudy (jdoe0001):

\(\large a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1)\quad \begin{cases} {\color{blue}{ d}}=common\ difference\\ {\color{brown}{ n}}=n^{th}\ term \end{cases}\)

OpenStudy (jdoe0001):

hmm \(\large a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1)\quad \begin{cases} a_1=1^{st}\ term\\ {\color{blue}{ d}}=common\ difference\\ {\color{brown}{ n}}=n^{th}\ term \end{cases}\)

OpenStudy (camerondoherty):

Maybe try each equation with one of the values for n and see which works?

OpenStudy (michele_laino):

@jdoe0001 please, you have to check in which among the formulae indicated, the numerical data are adapting

OpenStudy (michele_laino):

@camerondoherty yes!

OpenStudy (jdoe0001):

Michele_Laino I thought that is what's expected for CrazyCountryGirl to do

OpenStudy (michele_laino):

@jdoe Sorry, Sorry, I believed you got confused!

OpenStudy (jdoe0001):

hmmm I think I have a missing "d" but anyhow \(\large a_{\color{brown}{ n}}=a_1+({\color{brown}{ n}}-1){\color{blue}{ d}}\quad \begin{cases} a_1=1^{st}\ term\\ {\color{blue}{ d}}=common\ difference\\ {\color{brown}{ n}}=n^{th}\ term \end{cases}\)